The activation energy for the gas phase decomposition of difluoroperoxide is 72.4 kJ. F2O2F2 + O2 The rate constant at 238 K is 7.77×10-4 /s. The rate constant will be ___ /s at 279 K.
Answer #1

Solution:

The integrated Arhenious equation is written as:

ln (k2/k1) = Ea/R [T2-T1/T1T2]

Given that,

k1 = 7.77×10-4 sec-1 , T1 = 238 K , T2 = 279 K and Ea = 72.4 KJ = 72400 J, K2 =?

On putting the above values in equation, we get,

ln (k2/k1) = Ea/R [T2-T1/T1T2]

ln (k2/7.77×10-4 sec-1) =(72400 J /8.314 J K-1 mol-1) x [279 -238 /279 x 238]

ln (k2/7.77×10-4 sec-1) = 5.377

log (k2/7.77×10-4 sec-1) = 5.377/2.303 [ ln =2.303 log]

log (k2/7.77×10-4 sec-1) = 2.335

k2/7.77×10-4 sec-1 = antilog 2.335 = 216.27

k2 = 7.77×10-4 sec-1 x 216.27 = 1.68 x 10-1 = 0.168 sec-1